OpenStudy (anonymous):

What? Slope of y=-13/pi

OpenStudy (anonymous):

Since there is no x variable on the right hand side you are going to have a horizontal line at y = -13/pi So what would you say the slope is here?

OpenStudy (anonymous):

.........

OpenStudy (anonymous):

oh......y intercept will always be -4.13 right?

OpenStudy (anonymous):

Well at x = 1 it is y = -13/pi at x = 2 it is y = -13/pi If we use these two points (x, -13/pi) for (x1, y1) and (x, -13/pi) for (x2, y2) When we calculate slope: m = x2-x1/y2-y1 what do we get?

OpenStudy (anonymous):